OpenStudy (anonymous):

Find the value of the given trigonometric function. csc (-300

OpenStudy (ivettef365):

csc(x) = 1/sin x

OpenStudy (anonymous):

\[\frac{ -2 }{ \sqrt{3} }\]

OpenStudy (anonymous):

find \(-300\) or \(360-300=60\) on the unit circle on the last page of the attached cheat sheet the second coordinate of the point on the circle is sine take the reciprocal of that one

OpenStudy (anonymous):

\(-\frac{2}{\sqrt{3}}\) is not in fact correct, close though

OpenStudy (anonymous):

300 degree angle is in 4th quadrant and only cos and sec functions are positive in it

OpenStudy (anonymous):

oh sorry \[\frac{ 2 }{ \sqrt{3} }\]